OpenStudy (anonymous):

PLEASE HELP!!!!! Question 4: Part A: Factor 3x2y2 - 2xy2 - 8y2 . Show your work. (3 points) Part B: Factor x2 + 10x + 25. Show your work. (3 points) Part C: Factor x2 - 36. Show your work. (2 points) Part D: Do the polynomials in parts A, B, and C have a common factor? Explain your answer. (2 points) Answers: Part A: y^2(3x+4)(x-2) Part B: (x-5)^2 Part C: (x+6)(x-6) Part D: Yes the common factor is x^2

OpenStudy (anonymous):

www.mathway.com is a good website that will do all that work for you @CountryGurl15

OpenStudy (anonymous):

Okay really?

OpenStudy (anonymous):

you can factor out a y^2 in part a because all terms contain it

OpenStudy (anonymous):

Can you just tell if those r right or not please =)

OpenStudy (anonymous):

ohh sorry i didnt see the answers at the bottom

OpenStudy (anonymous):

part B is wrong

OpenStudy (anonymous):

yes those r the answers =)

OpenStudy (anonymous):

part b is wrong?

OpenStudy (ankit042):

Yup part B is incorrect (a+b)^2 = a^2+b^2 +2ab

OpenStudy (anonymous):

How would i fix part b?

OpenStudy (anonymous):

(x-5)^2 = x^2 + -10X + 25

OpenStudy (anonymous):

that is what your answers looks like factored

OpenStudy (anonymous):

oh okay so what would the correct answer be?

OpenStudy (ankit042):

check my previous message

OpenStudy (anonymous):

you need to make the 10 x positive by changing something.....

OpenStudy (anonymous):

adding 5?

OpenStudy (ankit042):

Nooooooo

OpenStudy (anonymous):

no inside the brackets

OpenStudy (anonymous):

(x-5)^2 to (x+5)^2 that's what i meant =)

OpenStudy (anonymous):

yeah nice job

OpenStudy (anonymous):

okay so that would make 10 positive ? =)

OpenStudy (anonymous):

yes, one way to check is to re distribute

OpenStudy (anonymous):

okay so it will be (x+5)^2 =x^2 +10x +25 ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

okay now what? =)

OpenStudy (anonymous):

so wait the factored version of x^2 +10 +25 is (x+5)^2 ?

OpenStudy (anonymous):

or (x+5)(x+5)

OpenStudy (anonymous):

okay that ^^ is the correct answer for part b?

OpenStudy (anonymous):

everything else looks good

OpenStudy (anonymous):

Awesome! thx so much for all the help! =)